# mpdframe frametitlebackgroundcolor =black!15, frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, skipabove =nermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm

### v2

@@ III-F Class of Functions Considered for Automated N-Version
Figure 5: Pure function example from the cryptography library libsodium, which meets the criteria for automated N-Version programming with Galápagos.
III-G Implementation
We implement the Diversification and Validation passes of Galápagos in Python, and make use of OpenAI’s GPT-4o API. The Harnessing pass is implemented in C++ in order to make use of LLVM’s code manipulation libraries. For formal equivalence checks, we rely on the alive-tv [30] tool that checks equivalence at the LLVM IR level.
@@ III-G Implementation
The code can be found in our open source repository at https://github.com/ASSERT-KTH/Galapagos.
IV Experimental Methodology
@@ 1. RQ1
    The variants generated by an LLM come with no guarantee. We aim to compute what percentage of these variants can be proven to perform the same task as the original function under diversification. To answer this question, we take a representative sample of functions from relevant C applications and libraries, and run them through Galápagos’ pipeline. Then, we analyze the distribution of these variants in terms of the different correctness filters of Galápagos’ validation passes presented in subsection III-D.
@@ 3. RQ3
    We evaluate the effectiveness of N-Version functions per our fault model, miscompilation bugs (see subsection III-A). We curate publicly documented, reproducible miscompilation bugs, and determine whether LLM-generated variants are useful in detecting and preventing them.
@@ 4. RQ4
IV-B Dataset of Functions
We select C functions from six open-source projects: alsa-lib, FFmpeg, libgcrypt, liboqs, libsodium, openssl. Those projects are all real-world, and have differing sizes, stages of maturity, and adoption.  Table I shows the selected projects, their versions, and related statistics. The number of commits for the selected projects’ repositories ranges between 1 458 and 111 662. Likewise, the number of C files and declared functions goes from 349 to 4 160, and 3 683 to 110 532 respectively.
@@ IV-B Dataset of Functions
TABLE I: List of real-world C projects used for experimental validation,. The versions marked with (*) represent the oldest named version before the commit used in this dataset.
From each of these projects, we select five functions that meet the following properties. (1) It is a pure function, and; (2) It does not call other functions. These properties ensure that the functions are supported by the semantic equivalence tool used in Galápagos’ verification step.
@@
TABLE II: List of 30 candidate real-world C functions for diversification.
Table II describes the set of functions we selected, which covers a diverse range of functionalities, e.g. arithmetic, type conversion, and cryptographic operations. It also shows each function’s size in lines of code, and the number of times that it is statically referenced within the project. The size of the functions by lines of code ranges from 2 to 34 with a median of 10. The number of times that each function is called within the same project ranges from 1 to 1 220 with a median of 2. To sum up, this dataset is exclusively composed of real-world C code that is used daily by millions of users.
@@
TABLE III: List of known miscompilation bugs selected for mitigation through Galápagos.
IV-C Methodology for RQ1: Assessing Functions Diversified by LLMs
@@
Importantly, we diversify the dataset of C functions (subsection IV-B) with two configurations same-language (C-to-C) and cross-language (C-to-Go). We produce 10 variants for each configuration, i.e. 100 variants per project, and 600 variants in total. Each variant is checked against all the filters described above, and the pass/fail occurrences are reported.
IV-D Methodology for RQ2: Assessing Variants Uniqueness
@@
Diversity at source code level can be removed through the transformations performed at different stages of compilation [35]. We measure how often this occurs, collecting the proportion of variants that preserve uniqueness after compilation with different optimization flags. We compare the unique SHA-256 hashes of the produced machine code. This experiment is performed on the same dataset of C functions as in RQ1 (IV-B). We perform this comparison only for variants that have been proven functionally equivalent previously by the filters applied in IV-C. We use each of four optimization flags -O0, -O1, -O2, -O3 to rebuild the function’s machine code, and compare the SHA256 hash of the produced binaries, to see whether some compilation flags undo specific variants (i.e. make the binary statically identical to the original).
@@
Then, we execute all statically diverse variants, to measure the diversity of observable behaviors among them. We define diverse observable behaviors per the CPU instruction trace. At runtime, we record all CPU instructions executed by the variant. To achieve this, we rely on Intel’s Pin instrumentation tool [36]. We then aggregate and compare the instructions executed by each variant in two ways. First, we compare the CPU instruction set used by the variants and compare it against the set of the original CPU instruction set. This is measured using the Jaccard similarity coefficient [37]. Second, we compare the amount of CPU instructions executed by the variants in terms of absolute numbers.
IV-E Methodology for RQ3: Validating the Ability to Mitigate Miscompilations
@@
As described in subsection III-A, our work specifically focuses on hardening against miscompilation errors produced by bugs in compilers, which is a very dangerous and tricky class of bugs. In RQ3, we consider miscompilation bugs from the Clang compiler. We assess the effectiveness of Galápagos in automatically generating N-Version programs which can mitigate these bugs. We look at all bugs reported in Table III, coming from different versions of Clang and optimization stages. All are unresolved in the repository’s issue tracker at the time of experimentation. For each bug, we create two test programs: one with a function containing a minimal reproducible example for said bug, and one with an N-Version implementation of the same function. We always run Galápagos with the minimal function as the reference broken run. We consider Galápagos to be successful if the N-Version implementation of the reference forces a crash instead of producing the incorrect output due to the miscompilation bug.
IV-F Methodology for RQ4: Measuring Performance Overhead
@@
V Experimental Results
In this section, we present and discuss the results of our experiments. In subsection V-A we look at the ability of Galápagos to generate variants that are equivalent to the reference (RQ1). In subsection V-B and subsection V-C, we assess the static and dynamic diversity of the variants, to answer RQ2. We conclude with subsection V-D, where we answer RQ3 with empirical evidence about the N-variants design’s ability to mitigate miscompilation errors.
V-A Validating Variants
@@
avg. 99.3% 96.0% 82.6% 32.3% 92.6% 76.3% 69.0% 24.3%
TABLE IV: Count of variant compliance to each validation filter: Compiled in isolation (Comp. Isol.), compiled within project (Comp. Proj.), test suite success (Test), and verified equal (Verif.).
To answer RQ1, we analyze the extent to which the generated variants pass the different validation filters. The numbers are presented in Table IV.
@@
The average rate of success decreases for test validation and equivalence verification to 82,6% (248/300) and 32,3% (93/300), respectively. Testing fails in 40 out of 288 remaining cases, either because of at least one failed test case or test suite timeout. An example from ffmpeg’s flac_get_max_frame_size is shown in 7, where constant addition to the count is not performed correctly. Finally, equivalence validation fails in 151 out of 248 remaining cases. Here, we uncover the cases where the LLM is able to synthesize variants that successfully run and behave similarly to the original, yet the complete semantics of the reference function are not correctly captured. Figure 8 shows an equivalence validation failure example for one of the variants of function fBlaMka. The figure shows a snippet of alive-tv’s that indicates that for inputs x and y of 131 074 and 51 111 936 respectively, the reference function outputs 13 398 943 041 538, while the variant outputs 7 235 012 610.
Figure 6: Build failure example: Function ‘__builtin_add_overflow‘ is not declared in the original project’s IR.
Figure 7: A logical error in variant (bottom) which causes test failure. The addition to count is not equivalent.
Figure 8: Formal equivalence validation failure: parameters x and y produce different outputs in the reference function and the variant.
@@
Figure 9: Build in isolation failure example. The code fails with an integer overflow error.
@@
{mdframed}
[style=mpdframe, frametitle=Answer to RQ1] Large language models successfully generate variants given a reference function. Out of 600 function variants, 170 were verifiably equivalent, 97 of which were created from same-language source code and 73 from a source code in a different language. This signals compelling applicability of creating automated cross-language N-Version functions with strong formal guarantees.
V-B Static Uniqueness
Table V shows the number of variants which compile to LLVM-unique IR code, and produce a unique machine code with different optimization flags. The column "Total eq." shows the total number of equivalent variants found, as displayed in Table IV. The column "IR" shows the number of these variants that are unique IR files. We obtain this number by comparing the "Total eq." variants’ hashes at the IR level. Subsequent columns, "-O0" to "-O3", show the number of unique machine code produced by the compiler with each optimization configuration.
For the same-language configuration, 91 of the 93 equivalent variants are different from each other at the IR level. These 91 variants are generated from 18 out of the 30 reference functions. The compilation with "-O0" preserves 80 unique variants. Meanwhile, the number of unique variants is reduced to 72 after "-O3" optimization, representing 74.22% of all equivalent variants. The optimizations observed in the functions are performed at the "-O1" level, and are concerned with inlining, loop invariant code motion, and global value numbering.
A key result for RQ2 is the fact that the introduced diversity is generally well-preserved across optimization stages: from all 127 unique variants at the IR level, a total of 93 (73.22%) are still unique after applying all optimization passes.
In total, at least 1 unique, equivalent variant was found for 24 out of 30 reference functions. Furthermore, we are able to find at least one behaviorally unique variant for each function for which we successfully synthesize at least 1 equivalent variant, . In the case of same-language diversification, the number of unique variants per function is greater than 2 in most cases (15/19). Cross-language diversification, on the other hand, does not perform as well, with only one unique variant for most functions (13/16). We hypothesize that the lower number of unique variants in the cross-language configuration is caused by two reasons: The LLM used in our experiment produces less diverse source code given our experimental parameters, or; the Go-to-IR compiler produces more consistent code across variants.
Project Function Name Same-language diversification (C) Cross-language diversification (C to Go) Totals
@@ V-B Static Uniqueness
Total 97 91 80 72 72 72 73 36 24 21 21 21 127 93
TABLE V: Variant uniqueness. Each column for both configurations shows how many unique variants are obtained at the IR level, and at the machine code level after various optimizations. Functions with no equivalent variants are omitted. The two rightmost columns display the total unique variants for both configurations, both at the IR level and at the machine code level, after building with the -O3 optimization flag.
V-C Dynamic Uniqueness
Now, we execute the diverse, equivalent variants generated by Galápagos. On average, same-language variants’ instruction sets have a Jaccard similarity coefficient of 0.805 with respect to the reference functions’ instruction sets; while the same metric for cross-language variants’ instruction sets is 0.693. In this measurement, the Jaccard similarity coefficient indicates overlap in observable internal behavior. Hence, cross-language variants are strictly better at maximizing diversity at runtime.
Figure 10 shows plots with CPU instruction counts of 6 of the dataset’s functions, one function per project. Each sub-figure shows 3 overlapping areas, aggregating the instruction counts of the original function, an equivalent same-language variant, and an equivalent cross-language variant. In detail, each figure is a radial plot, where the axes represent each of the CPU instructions executed by the function variants. Each colored area aggregates the instruction count of a different function implementation. This means that the less overlap in the figures, the more diverse the variants are in terms of the number and type of executed CPU instructions.
@@ V-C Dynamic Uniqueness
{mdframed}
[style=mpdframe, frametitle=Answer to RQ2] Our experiments show that Galápagos is able to create function variants that are diverse to a large extent. In total, for 166 unique variants across the 30 reference functions’ source code, 126 (75.9%) were still unique after all compilation and optimization passes. This is a very relevant finding, since it means that the approach is valid regardless of optimization requirements.
After executing the variants, we observe diverse internal behavior at runtime, evidenced by different sequences of CPU instructions. This finding is relevant, as it means that the introduced diversity at the machine code level goes beyond minor changes in the binary. It is aligned with the core assumption of N-Version programming, that the fault-tolerance increases with runtime diversity.
Figure 10: Dynamic uniqueness of variants. Each radial plot shows the count of CPU instructions executed by the original function, superimposed with the instruction counts of the same-language (C) and cross-language (Go) variants.
V-D Mitigating Miscompilations

### v3

@@ 3.6. Class of Functions Considered for Automated N-Version
Figure 5. Pure function example from the cryptography library libsodium, which meets the criteria for automated N-Version programming with Galápagos.
3.7. Implementation
We implement the Diversification and Validation passes of Galápagos in Python.
The LLM used in our experiments is OpenAI’s GPT-4o, because it was the best at the time of doing the experiments of this paper. The pipeline would support plugging other LLMs. It is left to future work to study the impact of varying the LLM on the diversification pipeline.
For formal equivalence checks, we rely on the alive-tv (lopes2021alive2, ) tool that checks equivalence at the LLVM IR level. alive-tv is an equivalence checker tied to the LLVM IR and its semantics. It operates by comparing two versions of a function: source and target. These versions must be well-formed LLVM IR and adhere to the same LLVM specification version. It verifies that the target function preserves the semantics of the source function, and enforces that the target function must be a refinement of the source function. This is, for all inputs where the source function has defined behavior, the target must produce the same observable result. The comparison is made by translating both variants to SMT and asks for logical equivalence. If the problem is satisfiable, it means that the solver has found a counter-example, that is an input such that both variants behave differently.
The Harnessing pass is implemented in C++ in order to make use of LLVM’s code manipulation libraries. It programmatically combines the LLVM IR version of the variants and calls the N-Version output comparison function.
@@ 3.7. Implementation
To provide maintanability and debugging in Galápagos, all outputs from each pass are saved to disk. This includes raw data from the LLM call, all variants’ source code and all resulting IR and machine code.
The implementation of Galápagos can be found in our open source repository at https://github.com/ASSERT-KTH/Galapagos.
4. Experimental Methodology
@@ 1. RQ1
    The variants generated by an LLM come with no guarantee. We aim to compute what percentage of these variants can be proven to perform the same task as the original function under diversification. To answer this question, we take a representative sample of functions from relevant C applications and libraries, and run them through Galápagos’ pipeline. Then, we analyze the distribution of these variants in terms of the different correctness filters of Galápagos’ validation passes presented in subsection 3.4.
@@ 3. RQ3
    We evaluate the effectiveness of N-Version functions per our fault model, miscompilation bugs (see subsection 3.1). We curate publicly documented, reproducible miscompilation bugs, and determine whether LLM-generated variants are useful in detecting and preventing them.
@@ 4. RQ4
4.2. Dataset of Functions
We select C functions from six open-source projects: alsa-lib, FFmpeg, libgcrypt, liboqs, libsodium, openssl. Those projects are all real-world, and have differing sizes, stages of maturity, and adoption.  Table 1 shows the selected projects, their versions, and related statistics. The number of commits for the selected projects’ repositories ranges between 1 458 and 111 662. Likewise, the number of C files and declared functions goes from 349 to 4 160, and 3 683 to 110 532 respectively.
@@ 4.2. Dataset of Functions
Table 1. List of real-world C projects used for experimental validation,. The versions marked with (*) represent the oldest named version before the commit used in this dataset.
From each of these projects, we select the first five functions that meet the following properties, analyzed sequentially. (1) It is a pure function, and; (2) It does not call other functions. These properties ensure that the functions are supported by the semantic equivalence tool used in Galápagos’ verification step.
@@
Table 2. List of 30 candidate real-world C functions for diversification. LOC "stands for lines of code". # Refs. stands for number of times that the function is statically referenced within the project.
Table 2 describes the set of functions we selected, which covers a diverse range of functionalities, e.g. arithmetic, type conversion, and cryptographic operations. It also shows each function’s size in lines of code (LOC), and the number of times that it is statically referenced within the project (#Refs). The size of the functions by lines of code ranges from 2 to 34 with a median of 10. The number of times that each function is called within the same project ranges from 1 to 1 220 with a median of 2. To sum up, this dataset is exclusively composed of real-world C code that is used daily by millions of users.
@@
Table 3. List of known miscompilation bugs selected for mitigation through Galápagos.
4.3. Methodology for RQ1: Assessing Functions Diversified by LLMs
@@
Importantly, we diversify the dataset of C functions (subsection 4.2) with two configurations same-language (C-to-C) and cross-language (C-to-Go). We produce 10 variants for each configuration, i.e. 100 variants per project, and 600 variants in total. Each variant is checked against all the filters described above, and the pass/fail occurrences are reported.
4.4. Methodology for RQ2: Assessing Variants Uniqueness
@@
Diversity at source code level can be removed through the transformations performed at different stages of compilation (cabrera-thesis, ). We measure how often this occurs, collecting the proportion of variants that preserve uniqueness after compilation with different optimization flags. We compare the unique SHA-256 hashes of the produced machine code, assuming that compilation is repeatable. This experiment is performed on the same dataset of C functions as in RQ1 (4.2). We perform this comparison only for variants that have been proven functionally equivalent previously by the filters applied in 4.3. We use each of four optimization flags -O0, -O1, -O2, -O3 to rebuild the function’s machine code, and compare the SHA256 hash of the produced binaries, to inspect whether some compilation flags undo specific variants (i.e. make the binary statically identical to the original).
@@
Then, we execute all statically diverse variants, to measure the diversity of observable behaviors among them. We define diverse observable behaviors per the CPU instruction trace. At runtime, we record all CPU instructions executed by the variant. To achieve this, we rely on Intel’s Pin instrumentation tool (intel-pin, ), since it allows us to record CPU instructions at function granularity, filtering out all other execution inherent to the used runtime, e.g. state initialization, garbage collection, etc. We then aggregate and compare the instructions executed by each variant in two ways. First, we compare the CPU instruction set used by the variants and compare it against the set of the original CPU instruction set. This is measured using the Jaccard similarity coefficient (jaccard, ). Second, we compare the amount of CPU instructions executed by the variants in terms of absolute numbers.
Finally, we compare the dynamic diversity of the LLM-generated variants against a baseline of rule-based variants generated with state-of-the-art tool Tigress (tigress, ). Tigress is a code transformation tool developed and maintained by Collberg and colleagues (collberg2012distributed, ) for diversification and obfuscation. We consider Tigress a suitable baseline approach since: (1) it works at the same level of abstraction; and (2) it aims to achieve behavioral diversity through program transformation, aligning with the objectives of N-Version programming. We apply the control-flow flattening transformation and use the resulting variants for comparison at runtime using the Jaccard similarity coefficient.
4.5. Methodology for RQ3: Validating the Ability to Mitigate Miscompilations
@@
As described in subsection 3.1, our work specifically focuses on hardening against miscompilation errors produced by bugs in compilers, which is a very dangerous and tricky class of bugs. In RQ3, we consider miscompilation bugs from the Clang compiler. We assess the effectiveness of Galápagos in automatically generating N-Version programs which can mitigate these bugs. We look at all bugs reported in Table 3, coming from different versions of Clang and optimization stages. All are unresolved in the repository’s issue tracker at the time of experimentation. For each bug, we create two test programs: one with a function containing a minimal reproducible example for said bug, and one with an N-Version implementation of the same function. We always run Galápagos with the minimal function as the reference broken run. We consider Galápagos to be successful if the N-Version implementation of the reference forces a crash instead of producing the incorrect output due to the miscompilation bug.
4.6. Methodology for RQ4: Measuring Performance Overhead
@@
5. Experimental Results
In this section, we present and discuss the results of our experiments. In subsection 5.1 we look at the ability of Galápagos to generate variants that are equivalent to the reference (RQ1). In subsection 5.2 and subsection 5.3, we assess the static and dynamic diversity of the variants, to answer RQ2. We conclude with subsection 5.4, where we answer RQ3 with empirical evidence about the resulting N-Version units’ ability to mitigate miscompilation errors.
5.1. Validating Variants
@@
avg. 99.3% 96.0% 82.6% 34.0% 92.6% 76.3% 69.0% 24.3%
Table 4. Count of variant compliance to each validation filter: Compiled in isolation (Comp. Isol.), compiled within project (Comp. Proj.), test suite success (Test), and verified equal (Verif.).
To answer RQ1, we analyze the extent to which the generated variants pass the different validation filters. The numbers are presented in Table 4.
@@
The average rate of success decreases for test validation and equivalence verification to 82,6% (248/300) and 34,3% (102/300), respectively. Testing fails in 40 out of 288 remaining cases, either because of at least one failed test case or test suite timeout. An example from ffmpeg’s flac_get_max_frame_size is shown in 7, where constant addition to the count is not performed correctly. Finally, equivalence validation fails in 151 out of 248 remaining cases. Here, we uncover the cases where the LLM is able to synthesize variants that successfully run and behave similarly to the original, yet the complete semantics of the reference function are not correctly captured. Figure 8 shows an equivalence validation failure example for one of the variants of function fBlaMka. The figure shows a snippet of alive-tv’s that indicates that for inputs x and y of 131 074 and 51 111 936 respectively, the reference function outputs 13 398 943 041 538, while the variant outputs 7 235 012 610.
Figure 6. Build failure example: Function ‘__builtin_add_overflow‘ is not declared in the original project’s IR.
Figure 7. A logical error in variant (bottom) which causes test failure. The addition to count is not equivalent.
Figure 8. Formal equivalence validation failure: parameters x and y produce different outputs in the reference function and the variant.
@@
Figure 9. Build in isolation failure example. The code fails with an integer overflow error.
@@
Large language models successfully generate variants given a reference function. Out of 600 function variants, 170 were verifiably equivalent, 97 of which were created from same-language source code and 73 from a source code in a different language. This signals compelling applicability of creating automated cross-language N-Version functions with strong formal guarantees.
5.2. Static Uniqueness
Project Function Name Same-language diversification (C) Cross-language diversification (C-to-Go) Totals
@@ 5.2. Static Uniqueness
Total 98 92 81 73 73 73 73 36 24 21 21 21 128 94
Table 5. Variant uniqueness. Each column for both configurations shows how many unique variants are obtained at the IR level, and at the machine code level after various optimizations. Functions with no equivalent variants are omitted. The two rightmost columns display the total unique variants for both configurations, both at the IR level and at the machine code level, after building with the -O3 optimization flag.
Table 5 shows the number of variants which compile to LLVM-unique IR code, and produce a unique machine code with different optimization flags. The column "Total eq." shows the total number of equivalent variants found, as displayed in Table 4. The column "IR" shows the number of these variants that are unique IR files. We obtain this number by comparing the "Total eq." variants’ hashes at the IR level. Subsequent columns, "-O0" to "-O3", show the number of unique machine code produced by the compiler with each optimization configuration.
For the same-language configuration, 92 of the 98 equivalent variants are different from each other at the IR level. These 92 variants are generated from 19 out of the 30 reference functions. The compilation with "-O0" preserves 81 unique variants. Meanwhile, the number of unique variants is reduced to 73 after "-O3" optimization, representing 74.48% of all equivalent variants. The optimizations observed in the functions are performed at the "-O1" level, and are concerned with inlining, loop invariant code motion, and global value numbering.
A key result for RQ2 is the fact that the introduced diversity is generally well-preserved across optimization stages: from all 128 unique variants at the IR level, a total of 94 (73.43%) are still unique after applying all optimization passes.
In total, at least 1 unique, equivalent variant was found for 24 out of 30 reference functions. Furthermore, we are able to find at least one behaviorally unique variant for each function for which we successfully synthesize at least 1 equivalent variant. In the case of same-language diversification, the number of unique variants per function is greater than 2 in most cases (15/19). Cross-language diversification, on the other hand, does not perform as well, with only one unique variant for most functions (13/16). We hypothesize that the lower number of unique variants in the cross-language configuration is caused by two reasons: The LLM used in our experiment produces less diverse source code given our experimental parameters, or; the Go-to-IR compiler produces more consistent code across variants.
5.3. Dynamic Uniqueness
Figure 10. Dynamic uniqueness of variants. Each radial plot shows the count of CPU instructions executed by the original function, superimposed with the instruction counts of the same-language (C) and cross-language (Go) variants.
Now, we execute the diverse, equivalent variants generated by Galápagos. Figure 10 shows plots with CPU instruction counts of 6 of the dataset’s functions, one function per project. Each sub-figure shows 3 overlapping areas, aggregating the instruction counts of the original function, an equivalent same-language variant, and an equivalent cross-language variant. In detail, each figure is a radial plot, where the axes represent each of the CPU instructions executed by the function variants. Each colored area aggregates the instruction count of a different function implementation. This means that the less overlap in the figures, the more diverse the variants are in terms of the number and type of executed CPU instructions.
@@ 5.3. Dynamic Uniqueness
Project Function Name C-to-C variants C-to-Go variants Tigress variant
Min Median Max Min Median Max
add 0.538 0.700 0.800 0.462 0.462 0.462 0.818
iec958_parity 0.750 0.750 0.833 0.500 0.636 0.636 0.733
alsa-lib val_seg 0.727 1.000 1.000 0.600
flac_get_max_frame_size 1.000 1.000 1.000 0.688
ffmpeg mix 1.000 1.000 1.000 0.714
barrett_reduce 0.900 0.900 0.900 0.769
ctz 0.462 0.462 0.643 0.444 0.538 0.667 0.750
int16_t_negative_mask 0.417 0.417 0.700 0.556 0.778 0.778 0.727
int16_t_nonzero_mask 0.357 0.455 0.700 0.308 0.308 0.308 0.727
libgcrypt montgomery_reduce 0.800 0.800 0.900 0.636 0.636 0.636 0.750
fpr_half 0.727 0.727 0.727 0.875 0.875 0.875 0.727
fpr_lt 0.412 0.412 0.412 0.750
int16_nonzero_mask 0.357 0.417 0.417 0.286 0.286 0.286 0.750
modp_montymul 0.750 0.750 1.000 0.769
liboqs modp_norm 1.000 1.000 1.000 0.727
libsodium fBlaMka 0.875 1.000 1.000 0.667
BitDeinterleave 0.364 0.500 0.500 0.571
BitInterleave 0.333 0.333 0.333 0.500 0.500 0.500 0.571
openssl _booth_recode_w5 1.000 1.000 1.000 1.000 1.000 1.000 0.786
Table 6. Jaccard coefficient comparison of executed x86 instructions to quantify runtime diversity introduced by Galápagos and Tigress’ flatten transform. For each function, we list the minimum, median, and maximum Jaccard coefficients for the C-to-C and C-to-Go variants produced by Galápagos, as well as the coefficient for the variant generated by Tigress. Lower coefficients indicate less overlap with the reference implementation and thus greater runtime diversity. Bold values indicate the greatest diversity (lowest coefficient) observed for each function. Missing entries (–) correspond to configurations where no equivalent variant could be produced.
Finally, Table 6 reports the overlap between the variants and their reference implementations, measured with the Jaccard coefficient over executed x86 instructions. Across 12 of the 19 functions, at least one variant produced by Galápagos attains a lower Jaccard coefficient than the variant produced by Tigress’ flatten transform. Overall, Galápagos’s variants exhibit lower coefficients on average than Tigress’, indicating less overlap and therefore greater runtime diversity. Moreover, the minimum observed Jaccard coefficient for Galápagos is 0.286, compared to 0.636 for Tigress’ flatten transform. Lower coefficients imply smaller overlaps in executed instructions, i.e., greater diversity at runtime.
Focusing on Galápagos’s variants, among the 11 functions where both same-language (C-to-C) and cross-language (C-to-Go) equivalents were found, 8 have their lowest coefficient coming from a cross-language variant. This suggests that cross-language variants are generally more effective at maximizing runtime diversity.
Our experiments show that Galápagos is able to create function variants that are diverse to a large extent, as measured through their persistence after compilation and optimization. In total, for 166 unique variants across the 30 reference functions’ source code, 126 (75.9%) remained unique after all compilation and optimization passes. This demonstrates that Galápagos’s approach to diversity is robust to different optimization requirements. We also compared Galápagos to a state-of-the-art rule-based diversification tool, Tigress. While Tigress is able to generate variants with runtime diversity, our results show that Galápagos generally achieves greater runtime diversity in the executed x86 instruction sequences. For 12/19 functions, the variants produced by Galápagos exhibit lower Jaccard coefficients relative to the reference implementation than those produced by Tigress, indicating less overlap and thus greater runtime diversity. This suggests that Galápagos’s translation-based approach can introduce a higher degree of behavioral diversity than traditional rule-based diversification transformations. In general for the generated variants, we observe diverse internal behavior at runtime, evidenced by different sequences of CPU instructions. This finding is relevant, as it means that the introduced diversity at the machine code level goes beyond minor changes in the binary. It is aligned with the core assumption of N-Version programming, that the fault-tolerance increases with runtime diversity.
5.4. Mitigating Miscompilations